OpenStudy (anonymous):

The length of a rectangle is 5 mm longer than its width. Its perimeter is more than 30 mm. Let w equal the width of the rectangle. Write an expression for the length in terms of the width. Use expressions for the length and width to write an inequality for the perimeter, on the basis of the given information. Solve the inequality, clearly indicating the width of the rectangle. Answer: helpp pleasee

OpenStudy (anonymous):

w=width w+5=length Perimeter = 2L+2s 30=2(w+5)+2w 30 = 2w+10+2w 30=4w+10 30-10=4w+10-10 20=4w 20/4=4w/4 5=w so the width is 5 and the length is 10

OpenStudy (anonymous):

thank you so much :)